OpenStudy (anonymous):

Using the equation 2x^2 + 10 = 12x: Find the vertex. Find the Discriminant. Graph the equation. Solve the equation by completing the square. Solve the equation using the quadratic formula.

OpenStudy (anonymous):

oh girl this is so long...will take forever to type

OpenStudy (anonymous):

i just need vertex & discriminant. the rest i can figure out myself @Best_Mathematician

OpenStudy (anonymous):

I think discriminant is \[b ^{2} - 4ac\]

OpenStudy (anonymous):

in this ques a = 2, b = 12, c = 10

OpenStudy (anonymous):

b = -12*

OpenStudy (anonymous):

vertex formula = -b/2a here a = 2, b = -12 and c = 10 so \[-\frac{ -12 }{ 2*2}\] =

OpenStudy (anonymous):

3

OpenStudy (anonymous):